10x + 7y = -120.5
-20x – 14y = 241

Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

10 x + 7 y = -120.5 .............1
-20 x -14 y = 241 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
20 x 14 y = -241
-20 x -14 y = 241
Add the two equations
0 x = 0
This is absurd
so the system has no solution. The lines are parallel

Answer by stanbon(75887)

10x + 7y = -120.5
-20x – 14y = 241
-----------------
Multiply thru the top equation by 2:
20x + 14y = -241
----
Add that to the bottom equation to get:
0 = 0
------
That means the two equations are really one equation.
There are an infinite number of solutions: (x ,(-10/7)x -(120.5/7))
+++++++++++++++++++++++=
Cheers,
Stan H.
=================================